OpenStudy (anonymous):

There are 3 white marbles and 7 blue marbles in a bag. Jamie will randomly pick two marbles out of the bag without replacing the first marble chosen. What is the probability of Jamie's picking 2 white marbles? A.) 1/15 B.) 1/18 C.) 1/20 D.) 1/7

OpenStudy (jango_in_dtown):

Let A denote the event of picking white marble at first and B denote the event of picking white marble again in the second.. Then P(A)=3/10 and P(B)=2/9. Required probability=P(AB)=P(A)P(B) since the events are independent =(3/10)(2/9)=1/15

OpenStudy (jango_in_dtown):

A

OpenStudy (jango_in_dtown):

@breezymeetee